OpenStudy (anonymous):

sin(x+45) = tanx, find x

OpenStudy (anonymous):

Your username is supporting your question.. :P

OpenStudy (anonymous):

Oh, from 18 minutes you have not got any help.. you wait has come to an end now.. :)

OpenStudy (anonymous):

First of all: \[\huge \color{green}{\textsf{Welcome To Openstudy...}}\]

OpenStudy (anonymous):

\[\sin(A+B) = \sin(A) \cos(B) + \cos(A)\sin(B)\]

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Umm..what??

OpenStudy (anonymous):

\[\implies \sin(x) \cos(45) + \cos(x) \sin(45) = \tan(x) \\ \implies \frac{1}{\sqrt{2}}(\sin(x) + \cos(x)) = \tan(x) \\ \implies Doing \; \; wrong..\]

OpenStudy (anonymous):

Nothing, just trying it.. :)

OpenStudy (anonymous):

The fact that wolfram alpha doesn't provide a form with respect to x (and the solutions it does provide are approximations) would suggest that this is impossible: http://www.wolframalpha.com/input/?i=sin(x%2B45)%3Dtan(x) I am approaching it the same as @waterineyes and I think we are both stuck in the same place. I would be very interested in seeing a solution though, if it is possible.

OpenStudy (anonymous):

Is the question from a textbook or did you make it up? :)

OpenStudy (anonymous):

I should also change my username now.. :P

OpenStudy (anonymous):

Oh its from some question paper There's no answer key... The question might be wrong...

ganeshie8 (ganeshie8):

what approximation methods are you allowed to use ?

OpenStudy (anonymous):

I'm not sure if other people noticed this but, by inspection, one solution is 45 degrees; this guy gives an awesome solution. https://answers.yahoo.com/question/index?qid=20130304110356AAQtCWL

ganeshie8 (ganeshie8):

nice :)

OpenStudy (anonymous):

x must be between 0 to 45 degrees.

OpenStudy (anonymous):

Ik x is 45 but idk how to get x as 45..

OpenStudy (anonymous):

Do you understand the solution given at Yahoo Answers or would you like an extra explanation? https://answers.yahoo.com/question/index?qid=20130304110356AAQtCWL

OpenStudy (anonymous):

And we aren't allowed to use any approximation methods..I'm still learning basic trig..

OpenStudy (anonymous):

Haha, 'basic' trig. Have you had a look at the Yahoo answers solution? Does it make sense? I would say that if you are learning "basic" trig then they would only be expecting you to look at the function and realise that 45 degrees would work.

OpenStudy (anonymous):

It kind of does, kind of doesn't.... But we can't just write 45..I mean there must be some way to solve it...

OpenStudy (anonymous):

There is but it is pretty difficult, the Yahoo answers solution is the only one that I can see. What grade are you in, in what country? (I am wondering what background knowledge you have)

OpenStudy (anonymous):

Well thanks anyways and I'm still in tenth so... I guess I'll just write 45.

OpenStudy (anonymous):

Nice work by Brian there.. :)

OpenStudy (anonymous):

Just by inspection we can see that if x = 45, then sin(45 + 45) = sin(90) = 1 and tan(45) = 1, so x = 45 degrees is one solution. (and in general (45 + n*360) degrees). But we should look to see if there are others. sin(x + 45) = sin(x)*cos(45) + cos(x)*sin(45) = (1/sqrt(2))*(sin(x) + cos(x)). So for this to equal tan(x) we require that sin(x) + cos(x) = sqrt(2)*tan(x) ----> sin(x) + cos(x) = sqrt(2)*sin(x)/cos(x) ----> sin(x)*cos(x) + cos^2(x) = sqrt(2)*sin(x) ----> cos^2(x) = (sqrt(2) - cos(x))*sin(x) ----> square both sides to get cos^4(x) = (2 - 2*sqrt(2)*cos(x) + cos^2(x))*sin^2(x) -----> cos^4(x) = (2 - 2*sqrt(2)*cos(x) + cos^2(x))*(1 - cos^2(x)) ----> cos^4(x) = 2 - 2*sqrt(2)*cos(x) - cos^2(x) + 2*sqrt(2)*cos^3(x) - cos^4(x) ----> 2*cos^4(x) - 2*sqrt(2)*cos^3(x) + cos^2(x) + 2*sqrt(2)*cos(x) - 2 = 0. Letting z = cos(x) the polynomial can be written as 2*z^4 - 2*sqrt(2)*z^3 + z^2 + 2*sqrt(2)*z - 2 = 0. We need numerical techniques to solve this.Using Newton's Method I get z = cos(x) = 0.707, (which gives the answer I guessed at initially), and z = cos(x) = -0.930, which gives us x = +/- 158.4 degrees + n*360 degrees. Testing this in the original equation, we find that only x = (158.4 + n*360) degrees works. So on [0, 360 degrees] the solutions are {45, 158.4 degrees}. (Note that the 158.4 degrees is only accurate to 1 decimal place.)